Minute of KAGRA commissioning meeting
- Date: 2018/5/8 11AM~12PM
- Attendance: Miyakawa, Yamamoto, Aso
- eZuce: Suzuki, Hirose, Yokozawa, Izumi, Ushiba
1. Discussion after operation
- IFO, ETMs: Izumi, Yamamoto
- Well done! Duty cycle is 80-90% in 1-5days, and 15-60% in 6-9days.
- IMC was quite stable.
- ETMY was dominant for stability of IFO in the last 4 days.
- Optical gain fluctuated 20~30% due to the contrast defect.
- BS coupling between L to P and Y was sometimes dominant for the contrast.
- Error of Sunup asymmetry was also dominant by optical gain.
- DAC of PR3 was dead. Restarting k1pr3 fixed this problem. We are not sure the cause.
- Timing of dc0 sometimes shifted. Restart daqd of k1dc0 fixed this problem. We are not sure the cause.
- Diaggui calculates float and sometimes coherence could not be measured.
- We switched VIS models to 2kHz to avoid glitches.
- Satellite box should be replaced to have more capacitance.
2. Discussion after phase1
- ETM leak was triggered 1week delay.
- Records of opening flange are very important.
- Manual for VIS and clean booth should be prepared.
- Mirror schedule should be checked frequently.
- Optimization of Each DOF of VIS takes 1 day, full VIS take 22 days.
- Hammaring test of ETMY showed that something hitting.
- Cable was well placed for ETMX. Fixing ETMY should be scheduled. (Ushiba)
- Heat link might be causes of fluctuation. It should be compared with ETMX.
- Needs ETMX cleaning. (Aso, Izumi, Hirose)
- TF of ETMY BF should be measured. (Izumi)
- Cross coupling of ETMY was large. (Izumi)
3. Phase 2
- Each stage and target
- Name of "Stage" should be replaced.
- 5/28 extend chief meeting for looking back phase 1.
- Next meeting: 5/15.